OpenStudy (anonymous):

how would i solve this in systems of equtions ) girl went to rodeo while there she saw 17 heads and 48 legs how many people did she see? hw manny calves?

OpenStudy (anonymous):

hello, before i help u with this equtions may i know what grade you r in?

OpenStudy (anonymous):

So we need variables to represent out unknowns: let x = Number of humans let y = Number of calves (mooo!) Unless we're in Roswell, humans and calves both have one head each. So this is our first equation. \[x+y=17\] We'll define y in terms of x (we will see why in a minute). \[y=17-x \Rightarrow y=-x+17\] (We could have done x in terms of y, it doesn't matter as we can substitute for either.)

OpenStudy (anonymous):

Now what is the relationship between Calves and Humans and the number of legs they have, which together is 48. ???? #Humans + ????? #Calves = 48 legs total. Well multiply the number of humans by 2 to get the number of human legs. Multiply the number of Calves by 4 to get the number of Calf legs. Our equation: \[2x+4y=48\] Now substitute using the definition for "y" from the first equation. Now what is the relationship between Calves and Humans and the number of legs they have, which together is 48. ???? #Humans + ????? #Calves = 48 legs total. Well multiply the number of humans by 2 to get the number of human legs. Multiply the number of Calves by 4 to get the number of Calf legs. Our equation: \[2x+4(-x+17)=48\] Distribute.. \[2x+(-4x)+68=48\] Go ahead and solve for x (the number of humans). You should be able to get the rest from here.